OpenStudy (anonymous):

Please help! My assignment is due a minute before midnight tonight. Find all the real zeros of the polynomial. Us the quadratic formula is necessary. P(x) = 2x^3 - 10x^2 +4x +24 Evaluate the expression and write the result in the form of a+bi. (Simplify completely) 8-7i (over) 1-8i (Last question is below)

OpenStudy (anonymous):

A polynomial P is given. P(x) = x^3 - 8 (a) Final all zeros of P, real and complex. (b) Factor P completely.

OpenStudy (anonymous):

we can do the last one fast

OpenStudy (anonymous):

\[x^3-8=(x-2)(x^2+2x+4)\] because that is how you factor the difference of two cubes evidently one zero is \(2\) the other two are found using the quadratic formula for \[x^2+2x+4=0\] either that, or what is even easier, completing the square

OpenStudy (anonymous):

you got that or you still need help otherwise we can move on and do the first two

OpenStudy (anonymous):

I got it but can you point out to me what the factored form is?

OpenStudy (anonymous):

sure once you get the zeros, which i believe are \(-1\pm\sqrt{3}i\) you factor as \[(x-2)(x-(1+\sqrt3 i))(x-(1-\sqrt3 i))\]

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

\[P(x) = 2x^3 - 10x^2 +4x +24\] solve \[P(x) = 2(x^3 - 5x^2 +2x +12)=0\]

OpenStudy (anonymous):

by some miracle one of the zeros is 3, so this factors as \[2 (x-3) (x^2-2 x-4) = 0\]

OpenStudy (anonymous):

Okay, I know 3 is one of the zeros.

OpenStudy (anonymous):

use the quadratic formula or complete the square to find the zeros of \[x^2-2x-4=0\]

OpenStudy (anonymous):

Sorry, internet running slow.

OpenStudy (anonymous):

completing the square is easiest, think you get \[x=1\pm\sqrt5\]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok so that one is done

OpenStudy (anonymous):

For the last question we have left I have done it before but two different people, one telling me it's right and another telling me it's wrong and I'm confused. Let me put up what I did.

OpenStudy (anonymous):

i have an idea, lets cheat

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

i can show you how to do it, most modern calculators can

OpenStudy (anonymous):

Ah, thank you.

OpenStudy (anonymous):

\[\frac{8-7i}{1-8i}\frac{1+8i}{1+8i}=\frac{(8-7i)(1+8i)}{1^2+8^2}\]

OpenStudy (anonymous):

almost midnight....

OpenStudy (anonymous):

Umm... Can we go back to the first question we did? Webassign isn't taking the factored form.

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

The P(x) = x^3 - 8

OpenStudy (anonymous):

if it does not like \[(x-2)(x-(1+\sqrt3 i))(x-(1-\sqrt3 i))\] try \[(x-2)(x-1-\sqrt3)(x-1+\sqrt3)\]

OpenStudy (anonymous):

and if it does not like that, then i am lost #$%^& webassign

OpenStudy (anonymous):

Alright. I thought webassign was the greatest thing at the beginning of the semester but I really hate with a Die Hard passion now.

OpenStudy (anonymous):

it is just frustrating when it wants one syntax and you write another

OpenStudy (anonymous):

Ah, I figured out what we did wrong.

OpenStudy (anonymous):

what ?

OpenStudy (anonymous):

We had \[(x-1-\sqrt{3}i)(x-1+\sqrt{3}i)\] instead of \[(x+1-\sqrt{3}i)(x+1+\sqrt{3}i)\]

OpenStudy (anonymous):

doh crap

OpenStudy (anonymous):

It's always the little things.

OpenStudy (anonymous):

And with 9 minutes to spare, haha.

OpenStudy (anonymous):

done? great time for me to reture

OpenStudy (anonymous):

retire

OpenStudy (anonymous):

Thank you for helping.

OpenStudy (anonymous):

yw